OpenStudy (anonymous):

The sum of 40 consecutive intergers is 100. What is the largest of these integers?

OpenStudy (anonymous):

hi let it an integer is x

OpenStudy (anonymous):

then x+x+1+x+2.........................x+40=100

hartnn (hartnn):

if the 1st integer is x then 40th integer = x+39 so, x+x+1+.....x+39 = 100

hartnn (hartnn):

now, notice that there will be 40 'x's, and the constants are 1+2+3+...38+39 to find, 1+2+3+...38+39 use, the formula for sum of first terms = \(n(n+1)/2\) here n=39. so, it'll be 40x+ 39*40/2 =100 or 40x + 20*39 = 100 now can u find 'x' from this equation ??